1/1+x^3 .dx Thank you.

1/1+x^3 .dx


Thank you.

Grade:

1 Answers

vikas askiitian expert
13 years ago

dear balaram

1+x3 = (1+x)(1+x2-x)

now integral becomes

I = dx/(1+x)(1+x2-x)

now use partial fraction 

1/(1+x)(1+x2-x) = A/(1+x) + (Bx+c)/(1+x2-x)

find A,B,C & then integrate it ...

approve my ans if u like it